# Fast Direct: Query-Efficient Online Black-box Guidance for Diffusion-model Target Generation

### v4

@@ 1 Introduction
Diffusion models have become the state-of-the-art generative model for image synthesis (Ho et al., 2020; Nichol & Dhariwal, 2021; Dhariwal & Nichol, 2021) and video synthesis (Ho et al., 2022), etc. Its remarkable success is due to its powerful capability in modeling the complex multi-mode high-dimensional data distributions.
One promising direction for utilizing the generative power of diffusion models is through target generation. This allows users to customize the generation process to meet specific downstream objectives, effectively extending the models’ capabilities beyond synthesis problems to engineering optimization and science discovery problems, such as image generation with human preferences, drug discovery (Corso et al., 2022; Guan et al., 2023) and material design (Vlassis & Sun, 2023; Giannone et al., 2023).
The pre-trained diffusion models often struggle to generate desired samples for these applications, especially when the target data lies out of the training data distribution. Therefore, the target generation often involves model fine-tuning or guidance techniques. Krishnamoorthy et al. (2023) proposes to train the diffusion model with re-weighted training loss, while Clark et al. (2023); Black et al. (2023); Fan et al. (2024); Yang et al. (2024) advocates for fine-tuning the parameters of the pre-trained model. As oppose to training-time approaches, Bansal et al. (2023) introduces an inference-time approach, which replaces the classifier in classifier guidance (Dhariwal & Nichol, 2021) with a differentiable objective function to achieve the downstream target. However, the requirement of the differentiable objective limits the practical usage of it for real applications with black-box feedback.
@@
Although the online black-box target generation tasks are important, the existing works are not suitable to address this task. Bansal et al. (2023); Krishnamoorthy et al. (2023); Lu et al. (2023) require training an offline guidance model with pre-collected data, while Bansal et al. (2023); Clark et al. (2023); Prabhudesai et al. (2023); He et al. (2023) require the objective function to be differentiable. Most recently, Black et al. (2023); Fan et al. (2024); Yang et al. (2024) can be employed for online black-box target generation, but they require online updates of the huge number of the parameters, which is both time-consuming and not query-efficient.
In Section 3.1, we first propose a novel guided noise sequence optimization (GNSO) technique to guide the diffusion model sampling process towards a given target. GNSO updates the noise sequence in a universal direction on the data manifold. Our GNSO is efficient: it enables fast adaptation (with around only 50 steps) from any initial generation towards the given input target. Moreover, GNSO is robust: empirically, even the input target image is noisy, it can still generate a clear image semantically similar to the target image. The GNSO itself cannot directly handle black-box target generation tasks, but it provides a backbone for further designing black-box algorithms.
In Section 3.2, based on our GNSO technique, we further propose a novel algorithm, Fast Direct, to tackle the online black-box target generation tasks in a query-efficient manner. Fast Direct builds a pseudo-target on the data manifold as the input for our GNSO, to guides the diffusion model at inference time. Notably, our Fast Direct is easy to implement and support any stochastic diffusion scheduler. Moreover, our Fast Direct provides a very flexible framework for extension. Any designs of update methods for the pseudo-target can serve as a plug-in for our Fast Direct.
@@
  • Fast Direct achieves significant query efficiency improvement in the real-world applications: high-resolution image target generation tasks and 3D-molecule target generation tasks. Additionally, we demonstrate its generalization capability for unseen prompts.
@@ 2.1 Guided Generation
Dhariwal & Nichol (2021) proposed classifier guidance. However, it requires a guidance model that trained on noisy images with different noise scales, which generally not readily available and often requires training from scratch for each domain. Chung et al. (2022); Bansal et al. (2023); He et al. (2023) extend the classifier guidance by using a differentiable objective function that is defined only for clean data. Instead of using noisy images, the predicted clean images at each sampling steps are used as the input for the guidance objective function. In this way, the guidance process can operate on the clean image space. While the predicted clean image is naturally imperfect, empirically it still provide informative feedback to guide image generation (Bansal et al., 2023).
However, this approximation can harm image quality. Bansal et al. (2023) proposed universal guidance that comprises of backward guidance followed by a self-recurrence step to preserve image quality. On the other hand, He et al. (2023) leverages the differentiability of a well-trained auto-encoder to project the image to the data manifold and thus preserve image quality during the guidance process. Instead of guiding the sampling trajectory, Karunratanakul et al. (2024); Eyring et al. (2024) treat the diffusion process as a black-box and only optimize the initial (prior) noise.
While the aforementioned approaches assume a differentiable objective function, Lu et al. (2023) tackles black-box objective $f$ by learning a differentiable proxy neural network $h$ to match their gradients, i.e., $\nabla f\approx\nabla h$. Li et al. (2024) eliminates the need for a differentiable proxy model by employing importance sampling weighted by the objective values during the sampling process. Most recently, DNO (Tang et al., 2024) proposes optimize the diffusion noise sequence by using ZO-SGD (Nesterov & Spokoiny, 2017) to tackle black-box objective function. However, it runs at the instance level; namely, each run only produces one image.
@@ 2.2 Diffusion-model Fine-tuning
Diffusion-model fine-tuning refer to the technique that updating the pre-trained diffusion-model parameters to improve its performance on a specific use case. In this section, we review the fine-tuning methods that support online learning of the black-box objective function.
Black et al. (2023); Fan et al. (2024) formulate the diffusion fine-tuning problem as a reinforcement learning (RL) problem within Markov Decision Processes (MDPs), and proposes an iterative algorithm to fine-tune diffusion model by using Proximal Policy Optimization (PPO) (Schulman et al., 2017; Uehara et al., 2024) loss function. Fan et al. (2024) integrates the PPO with a KL regularization to prevent the fine-tuned model deviated too much from the pre-trained model.
On the other hand, Yang et al. (2024) does not requires an absolute objective values, instead it uses the relative reward on pair of samples by integrating DPO (Direct Preference Optimization) (Rafailov et al., 2024), a technique for fine-tuning large language models, into diffusion model.
Fine-tuning diffusion model requires large amount of GPU memory. Existing works mitigates the memory consumption by using LoRA (Low-Rank Adaptation) (Hu et al., 2021) technique, and only fine-tune parameters of the attention blocks in UNet. We categorize the related works based on whether they support the online and black-box objective tasks in Appendix F Table 2.
@@ 12             $\hat{\bm{\epsilon}}_{k}\leftarrow\bm{\epsilon}_{k}+\alpha(\bm{x}^{*}-\bm{x}_{K})$
Figure 1: Demonstration of guided generation for a given target by Algorithm 1. Column 2 (Update Direction) indicates the update term of Algorithm 1 Line 8. Row 1 to 3 analyze how different update directions can affect the generated images. Row 4 and 5 show that by using the update direction of $\bm{x}^{*}-\bm{x}_{K}$, diffusion model is able to generate visually satisfied images even when the target image is noisy. The noisy target image ($\bm{x}^{*}$) of the row 4 is obtained by clean image added with noise $\mathcal{N}(\bm{0},\bm{I})$, and row 5 is added with noise $\mathcal{N}(\bm{0},9\times\bm{I})$.
@@
Question 1: Given an input target $\bm{x}^{*}$, can we guide a pre-trained diffusion model at inference time to generate the target data in an efficient and robust manner?
To answer this question, we need to develop an algorithm to address two challenges simultaneously: (1) Guidance Efficiency: the algorithm can use only a few iteration steps of guidance update to generate target data $\bm{x}^{*}$. (2) Robust to noisy target: given a noisy target (e.g., a noise perturbed image), the algorithm can generate meaningful target data (e.g., a clear image).
We propose a novel guided noise sequence optimization method to address the above two challenges, which is presented in Algorithm 1. Our method optimizes the noise sequence by updating noise along direction on data manifold. Specifically, Algorithm 1 consists of two procedures: data generation procedure and noise update procedure. In data generation procedure, Line 4 - Line 6 in Algorithm 1 employ current noise sequence $\{\bm{\epsilon}_{0},\cdots,\bm{\epsilon}_{K}\}$ to generate a data $\bm{x}_{K}$. In noise update procedure, Line 7 - Line 9 in Algorithm 1, we update each noise $\bm{\epsilon}_{k}$ by moving along a universal direction $\bm{x}^{*}-\bm{x}_{K}$. Then, Algorithm 1 continues to employ the updated noise to generate a new data $\bm{x}_{K}$. By looping the data generation procedure and noise update procedure, Algorithm 1 can generate a clear and meaningful target.
@@
We empirically evaluate Algorithm 1 with different update directions: $\vec{d}=\bm{x}^{*}-\bm{x}_{K}$ and $\vec{d}=\bm{x}^{*}-\bm{x}_{k}$. Additionally, we consider the approximation $\bm{x}_{K}\approx\hat{\bm{x}}_{k}$ [^1] The $\hat{\bm{x}}_{k}$ is the ”predicted clean image” at step $k$, see the Eq.12 of DDIM (Song et al., 2020). Note that it is undefined for initial step, so we simply set $\hat{\bm{x}}_{0}\coloneqq\hat{\bm{x}}_{1}$.. Lastly, we set the noisy target to test the robustness. We use the stable-diffusion model as the pre-trained model. The stepsize $\alpha$ is set to $\alpha=2\times 10^{-3}$. We present the generated images at every 10-iteration for each cases in Figure 1.
@@
Moreover, we present the updating direction $\bm{x}^{*}-\bm{x}_{K^{\prime}}$ for $K^{\prime}\in\{K,K/2,K/4,K/8\}$ in Appendix C. We observe that the generated image quality decreases as the $\bm{x}_{K^{\prime}}$ becomes more noisy.
@@
To answer this question, we need to address two challenges: (1) Black-box challenge and (2) Online guidance challenge. The black-box challenge means that we cannot access the gradient of the objective. The online guidance challenge means we don’t have a prior dataset to train a surrogate (classifier) for guidance; we can only access the black-box objective through query feedback online. Usually, the query evaluation is expensive. Thus, the query efficiency is critical.
We start addressing the above two challenges based on our target guidance generation Algorithm 1. Although Algorithm 1 itself cannot handle black-box guidance tasks because it needs input a target $\bm{x}^{*}$, it provides a basis for us to design query efficient online black-box guidance methods. To be specific, Algorithm 1 enables us to generate a target even updating with a noisy direction $\bm{x}^{*}-\bm{x}_{K}$. This property is important for black-box guidance tasks because we can use a noisy estimation of the gradient of the black-box objective to guide the generation.
@@ 18                         $\hat{\bm{\epsilon}}_{k}\leftarrow\bm{\epsilon}_{k}+\alpha(\hat{\bm{x}}^{*}-\bm{x}_{K})$
Remark: When setting the pseudo target $\hat{\bm{x}}^{*}$ as Eq.(2), we know the pseudo target $\hat{\bm{x}}^{*}$ lies on the low-dimensional subspace spanned by data according to Proposition 1. Usually, the gradient descent update in Eq.(2) will lead to a new pseudo target $\hat{\bm{x}}^{*}$ with a lower score. In addition, based on the empirical observation from Algorithm 1, we know even the pseudo target $\hat{\bm{x}}^{*}$ is noisy, Algorithm 1 can still generate similar data that is meaningful lies on the data manifold. These two properties enable Algorithm 2 to generate data with lower and lower scores along the data manifold.
@@
Set pseudo target $\hat{\bm{x}}^{*}$ as Eq.(4) is an empirical approximation of the optimal target $\bm{x}^{*}=\operatorname*{arg\,min}_{\bm{x}\in\mathcal{X}}f(\bm{x})$. In Algorithm 2, at each iteration w.r.t. the batch query $i$, we employ the generated data that achieves the current best score as the pseudo target $\hat{\bm{x}}^{*}$ to guide the generation process. Intuitively, given this pseudo target $\hat{\bm{x}}^{*}$, Algorithm 2 will generate a batch of data that is similar to $\hat{\bm{x}}^{*}$ on the low-dimensional data manifold, which is promising to further improve the score. By looping this procedure, we will get data generation with lower and lower scores.
@@ 4 Experiments
In this section, we evaluate our algorithm in two domains, images and molecules, and compare it against four baseline methods: DDPO (Black et al., 2023), DPOK (Fan et al., 2024), D3PO (Yang et al., 2024), DNO (Tang et al., 2024). Moreover, we further evaluate our algorithm in compressibility, incompressibility, and aesthetic quality tasks in Appendix A.

### v5

@@ 1 Introduction
Diffusion models have become the state-of-the-art generative model for image synthesis (Ho et al., 2020; Nichol & Dhariwal, 2021; Dhariwal & Nichol, 2021) and video synthesis (Ho et al., 2022), etc. Its remarkable success is due to its powerful capability in modeling complex multi-mode high-dimensional data distributions.
One promising direction for utilizing the generative power of diffusion models is through target generation, which allows users to customize the generation process to meet specific downstream objectives, effectively extending the models’ capabilities beyond synthesis problems to engineering optimization and science discovery problems, such as image generation with human preferences, drug discovery (Corso et al., 2022; Guan et al., 2023) and material design (Vlassis & Sun, 2023; Giannone et al., 2023).
The pre-trained diffusion models often struggle to generate desired samples for these applications, especially when the target data lies outside the training data distribution. Therefore, the target generation often involves model fine-tuning or guidance techniques. Krishnamoorthy et al. (2023) proposes to train the diffusion model with re-weighted training loss, while Clark et al. (2023); Black et al. (2023); Fan et al. (2024); Yang et al. (2024) advocates for fine-tuning the parameters of the pre-trained model. As opposed to training-time approaches, Bansal et al. (2023) introduces an inference-time approach, which replaces the classifier in classifier guidance (Dhariwal & Nichol, 2021) with a differentiable objective function to achieve the downstream target. However, the requirement of the differentiable objective limits its practical usage for real applications with black-box feedback.
@@
Although the online black-box target generation tasks are important, the existing works are not suitable to address this task. Bansal et al. (2023); Krishnamoorthy et al. (2023); Lu et al. (2023) require training an offline guidance model with pre-collected data, while Bansal et al. (2023); Clark et al. (2023); Prabhudesai et al. (2023); He et al. (2023) require the objective function to be differentiable. Most recently, Black et al. (2023); Fan et al. (2024); Yang et al. (2024) can be employed for online black-box target generation, but they require online updates of the huge number of parameters, which is both time-consuming and not query-efficient.
In Section 3.1, we first propose a novel guided noise sequence optimization (GNSO) technique to guide the diffusion model sampling process towards a given target. GNSO updates the noise sequence in a universal direction on the data manifold. Our GNSO is efficient: it enables fast adaptation (with around only 50 steps) from any initial generation towards the given input target. Moreover, GNSO is robust: empirically, even if the input target image is noisy, it can still generate a clear image semantically similar to the target image. The GNSO itself cannot directly handle black-box target generation tasks, but it provides a backbone for further designing black-box algorithms.
In Section 3.2, based on our GNSO technique, we further propose a novel algorithm, Fast Direct, to tackle the online black-box target generation tasks in a query-efficient manner. Fast Direct builds a pseudo-target on the data manifold as the input for our GNSO, to guide the diffusion model at inference time. Notably, our Fast Direct is easy to implement and supports any stochastic diffusion scheduler. Moreover, our Fast Direct provides a very flexible framework for extension. Any designs of update methods for the pseudo-target can serve as a plug-in for our Fast Direct.
@@
  • Fast Direct achieves significant query efficiency improvement in real-world applications: high-resolution image target generation tasks and 3D-molecule target generation tasks. Additionally, we demonstrate its generalization capability for unseen prompts.
@@ 2.1 Guided Generation
Dhariwal & Nichol (2021) proposed classifier guidance. However, it requires a guidance model trained on noisy images with different noise scales, which is generally not readily available and often requires training from scratch for each domain. Chung et al. (2022); Bansal et al. (2023); He et al. (2023) extend the classifier guidance by using a differentiable objective function defined only for clean data. Instead of using noisy images, the predicted clean images at each sampling step are used as the input for the guidance objective function. In this way, the guidance process can operate on the clean image space. While the predicted clean image is naturally imperfect, empirically it still provides informative feedback to guide image generation (Bansal et al., 2023).
However, this approximation can harm image quality. Bansal et al. (2023) proposed universal guidance that comprises backward guidance followed by a self-recurrence step to preserve image quality. On the other hand, He et al. (2023) leverages the differentiability of a well-trained auto-encoder to project the image to the data manifold and thus preserve image quality during the guidance process. Instead of guiding the sampling trajectory, Karunratanakul et al. (2024); Eyring et al. (2024) treat the diffusion process as a black-box and only optimize the initial (prior) noise.
While the aforementioned approaches assume a differentiable objective function, Lu et al. (2023) tackles black-box objective $f$ by learning a differentiable proxy neural network $h$ to match their gradients, i.e., $\nabla f\approx\nabla h$. Li et al. (2024) eliminates the need for a differentiable proxy model by employing importance sampling weighted by the objective values during the sampling process. Most recently, DNO (Tang et al., 2024) proposes optimizing the diffusion noise sequence by using ZO-SGD (Nesterov & Spokoiny, 2017) to tackle the black-box objective function. However, it runs at the instance level; namely, each run only produces one image.
@@ 2.2 Diffusion-model Fine-tuning
Diffusion-model fine-tuning refers to the technique that updating the pre-trained diffusion-model parameters to improve its performance on a specific use case. In this section, we review the fine-tuning methods that support online learning of the black-box objective function.
Black et al. (2023); Fan et al. (2024) formulate the diffusion fine-tuning problem as a reinforcement learning (RL) problem within Markov Decision Processes (MDPs), and propose an iterative algorithm to fine-tune the diffusion model by using Proximal Policy Optimization (PPO) (Schulman et al., 2017; Uehara et al., 2024) loss function. Fan et al. (2024) integrates the PPO with a KL regularization to prevent the fine-tuned model from deviating too much from the pre-trained model.
On the other hand, Yang et al. (2024) does not require an absolute objective value; instead, it uses the relative reward on a pair of samples by integrating DPO (Direct Preference Optimization) (Rafailov et al., 2024), a technique for fine-tuning large language models, into the diffusion model.
Fine-tuning the diffusion model requires a large amount of GPU memory. Existing work mitigates memory consumption by using the LoRA technique (Low-Rank Adaptation) (Hu et al., 2021) and only fine-tunes parameters of the attention blocks in UNet. We categorize the related works based on whether they support the online and black-box objective tasks in Appendix F Table 2.
@@ 12             $\hat{\bm{\epsilon}}_{k}\leftarrow\bm{\epsilon}_{k}+\alpha(\bm{x}^{*}-\bm{x}_{K})$
Figure 1: Demonstration of guided generation for a given target by Algorithm 1. Column 2 (Update Direction) indicates the update term of Algorithm 1, Line 8. Rows 1 to 3 analyze how different update directions can affect the generated images. Rows 4 and 5 show that by using the update direction of $\bm{x}^{*}-\bm{x}_{K}$, the diffusion model can generate visually satisfying images even when the target image is noisy. The noisy target image ($\bm{x}^{*}$) of the row 4 is obtained by clean image added with noise $\mathcal{N}(\bm{0},\bm{I})$, and row 5 is added with noise $\mathcal{N}(\bm{0},9\times\bm{I})$.
@@
Question 1: Given an input target $\bm{x}^{*}$, can we guide a pre-trained diffusion model at inference time to generate the target data efficiently and robustly?
To answer this question, we need to develop an algorithm to address two challenges simultaneously: (1) Guidance Efficiency: the algorithm can generate the target data $\bm{x}^{*}$ by using only a few guidance update steps. (2) Robust to noisy target: given a noisy target (e.g., a noise-perturbed image), the algorithm can generate meaningful target data (e.g., a clear image).
We propose a novel guided noise sequence optimization method to address the above two challenges, which is presented in Algorithm 1. Our method optimizes the noise sequence by updating the noise along the direction on the data manifold. Specifically, Algorithm 1 consists of two procedures: a data generation procedure and a noise update procedure. In the data generation procedure, Line 4 - Line 6 in Algorithm 1 employ current noise sequence $\{\bm{\epsilon}_{0},\cdots,\bm{\epsilon}_{K}\}$ to generate a data $\bm{x}_{K}$. In the noise update procedure, Lines 7 - 9 in Algorithm 1, we update each noise $\bm{\epsilon}_{k}$ by moving along a universal direction $\bm{x}^{*}-\bm{x}_{K}$. Then, Algorithm 1 continues to employ the updated noise to generate a new data $\bm{x}_{K}$. By looping the data generation procedure and noise update procedure, Algorithm 1 can generate a clear and meaningful target.
@@
We empirically evaluate Algorithm 1 with different update directions: $\vec{d}=\bm{x}^{*}-\bm{x}_{K}$ and $\vec{d}=\bm{x}^{*}-\bm{x}_{k}$. Additionally, we consider the approximation $\bm{x}_{K}\approx\hat{\bm{x}}_{k}$ [^1] The $\hat{\bm{x}}_{k}$ is the ”predicted clean image” at step $k$, see the Eq.12 of DDIM (Song et al., 2020). Note that it is undefined for the initial step, so we simply set $\hat{\bm{x}}_{0}\coloneqq\hat{\bm{x}}_{1}$.. Lastly, we set the noisy target to test the robustness. We use the stable-diffusion model as the pre-trained model. The stepsize $\alpha$ is set to $\alpha=2\times 10^{-3}$. We present the generated images every 10 iterations for each case in Figure 1.
@@
Moreover, we present the updating direction $\bm{x}^{*}-\bm{x}_{K^{\prime}}$ for $K^{\prime}\in\{K,K/2,K/4,K/8\}$ in Appendix C. We observe that the generated image quality decreases as the $\bm{x}_{K^{\prime}}$ becomes noisier.
@@
To answer this question, we need to address two challenges: (1) the Black-box challenge and (2) the Online guidance challenge. The black-box challenge means that we cannot access the gradient of the objective. The online guidance challenge means we don’t have a prior dataset to train a surrogate (classifier) for guidance; we can only access the black-box objective through query feedback online. Usually, the query evaluation is expensive. Thus, query efficiency is critical.
We start addressing the above two challenges based on our target guidance generation Algorithm 1. Although Algorithm 1 itself cannot handle black-box guidance tasks because it needs input a target $\bm{x}^{*}$, it provides a basis for us to design query-efficient online black-box guidance methods. To be specific, Algorithm 1 enables us to generate a target even updating with a noisy direction $\bm{x}^{*}-\bm{x}_{K}$. This property is important for black-box guidance tasks because we can use a noisy estimation of the gradient of the black-box objective to guide the generation.
@@ 18                         $\hat{\bm{\epsilon}}_{k}\leftarrow\bm{\epsilon}_{k}+\alpha(\hat{\bm{x}}^{*}-\bm{x}_{K})$
Remark: When setting the pseudo target $\hat{\bm{x}}^{*}$ as Eq.(2), we know the pseudo target $\hat{\bm{x}}^{*}$ lies on the low-dimensional subspace spanned by data according to Proposition 1. Usually, the gradient descent update in Eq.(2) will lead to a new pseudo target $\hat{\bm{x}}^{*}$ with a lower score. In addition, based on the empirical observation from Algorithm 1, we know that even the pseudo target $\hat{\bm{x}}^{*}$ is noisy, Algorithm 1 can still generate similar data that is meaningful and lies on the data manifold. These two properties enable Algorithm 2 to generate data with lower and lower scores along the data manifold.
@@
Set pseudo target $\hat{\bm{x}}^{*}$ as Eq.(4) is an empirical approximation of the optimal target $\bm{x}^{*}=\operatorname*{arg\,min}_{\bm{x}\in\mathcal{X}}f(\bm{x})$. In Algorithm 2, at each iteration w.r.t. the batch query $i$, we employ the generated data that achieves the current best score as the pseudo target $\hat{\bm{x}}^{*}$ to guide the generation process. Intuitively, given this pseudo target $\hat{\bm{x}}^{*}$, Algorithm 2 will generate a batch of data that is similar to $\hat{\bm{x}}^{*}$ on the low-dimensional data manifold, which is promising to further improve the score. By iteratively repeating this procedure, the generated data progressively achieves lower objective scores.
@@ 4 Experiments
In this section, we evaluate our algorithm in two domains, images and molecules, and compare it against four baseline methods: DDPO (Black et al., 2023), DPOK (Fan et al., 2024), D3PO (Yang et al., 2024), and DNO (Tang et al., 2024). Moreover, we further evaluate our algorithm in compressibility, incompressibility, and aesthetic quality tasks in Appendix A.